OpenStudy (anonymous):

Which equation is not an identity? tan(–t) = –tan t tan t(csc t + cot t) = sec t + 1 sin tcot t = cos t cos2t − sin2t = –1

OpenStudy (anonymous):

@Shalante @oldrin.bataku

OpenStudy (anonymous):

i feel like it is d but in like 50% sure

OpenStudy (anonymous):

@UsukiDoll

OpenStudy (usukidoll):

thinking...

OpenStudy (usukidoll):

I've seen \[\cos^2t-\sin^2t \] before. I think it was one of the identities for cos2t

OpenStudy (anonymous):

okay so d is right?

OpenStudy (usukidoll):

yeah... there's also another identity which is \[\large \cos^2t+\sin^2t=1 \] but rearranging them a bit gives me \[\large \cos^2t=1-\sin^2t \] \[\large \sin^2t=1-\cos^2t \]

OpenStudy (anonymous):

d was right lol this stuff is so confusing

OpenStudy (usukidoll):



OpenStudy (usukidoll):

one big huge trig list........and parts of it exist beyond trig.

OpenStudy (anonymous):

okay i posted another question do you think you could help me on that one as well?